Celsius temperature readings can be converted to fahrenheit readings using the formular F = 9 over 5 C+ 32. What is the fahrenheit tempature that corresponds to each of the following celsius temperatures:
15

Answer by fractalier(2101)

The conversion formula is
F = (9/5)C + 32
F = (9/5)(15) + 32
F = 27 + 32
F = 59